OpenStudy (anonymous):

What is the slope between the two points (4, 14) and (2, 4)?

OpenStudy (anonymous):

slope = rise over run (y2-y1)/(x2-x1) (14-4)7(4-2)=10/2=5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob